OpenStudy (anonymous):

A 16-foot board is cut into shelves, each 3 1/3 feet in length. What is the length of the remaining piece? A)2^ 2/3 feet B)1 ^2/3 feet C) 2^ 1/3 feet D)1^ 1/3 feet

OpenStudy (anonymous):

 \[x = 16 - 4 \times 3\frac{1}{3}\]